# Editing Personality for LLMs

### v1

@@ Abstract
This paper introduces an innovative task focused on editing the personality traits of Large Language Models (LLMs). This task seeks to adjust the models’ responses to opinion-related questions on specified topics since an individual’s personality often manifests in the form of their expressed opinions, thereby showcasing different personality traits. Specifically, we construct a new benchmark dataset PersonalityEdit to address this task. Drawing on the theory in Social Psychology (Goldberg, 1990), we isolate three representative traits, namely Neuroticism, Extraversion, and Agreeableness, as the foundation for our benchmark. We then gather data using GPT-4, generating responses that not only align with a specified topic but also embody the targeted personality trait. We conduct comprehensive experiments involving various baselines and discuss the representation of personality behavior in LLMs. Our intriguing findings uncover potential challenges of the proposed task, illustrating several remaining issues. We anticipate that our work can provide the NLP community with insights [^1] Code and datasets will be released at https://github.com/zjunlp/EasyEdit..
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable strides in modeling language distributions and excelling in a wide array of NLP tasks (OpenAI, 2023a; Yao et al., 2023a; Zhao et al., 2023; Yin et al., 2023). More recent studies (Park et al., 2023; Akata et al., 2023; Wang et al., 2023d; b; Zhou et al., 2023; Xi et al., 2023) have further expanded our understanding of LLMs in role-playing scenarios, which effectively serve as a rich array of agents, embodying a multitude of potential char dacters within an expansive multiverse (Shanahan et al., 2023).
Unlike LLMs, humans exhibit distinct personalities, and each person has a certain degree of personality in their response to events and actions (Goldberg, 1981). This leads to the research question: Can we edit the personality for LLMs? Note that editing personality for LLMs can: 1) precisely customize and edit the behavioral expressions of LLMs; 2) personalize LLMs to meet the needs of different users and scenarios; 3) help analyze the ethics and safety of LLMs.
To address this need, we take the first step to construct PersonalityEdit, a new benchmark for a comprehensive evaluation of editing personality for LLMs. This inspiration is drawn from the big-five factor structure in Social Psychology (Goldberg, 1990). Specifically, as shown in Figure 1, we focus on three of the Big Five personality traits: Neuroticism, Extraversion, and Agreeableness, because Extraversion and Neuroticism are more comprehensible in terms of their foundational processes (DeYoung et al., 2010), coupled with the distinctive nature of Agreeableness compared to the other traits. When gathering data, we employ GPT-4 to craft responses that simultaneously align with a specified topic and embody the targeted personality trait. For quality control, we utilize automatic methods supplemented with human verification to filter the data, which yield 14,400 instances for training, 1,800 for development, and 1,800 for testing.
We conduct a comprehensive evaluation with multiple representative model editing methods, utilizing two kinds of mainstream LLMs within the context of the proposed benchmark. Empirically, previous baselines can implement personality editing to some extent, but the effect is still barely satisfactory, indicating the potential difficulty of this task. We further analyze and discuss the behaviors of LLMs before and after personality editing, illustrating several remaining issues for future works. We summarize the major contributions of this work as follows:
@@
  • We pioneer the approach of employing GPT-4 for topic-constrained and personality trait-guided data generation. Then we implement automated methods as well as meticulous human verification to ensure the utmost quality control.
  • For thorough experiments, we propose several metrics to evaluate personality traits in the generated text. We analyze different baselines, revealing that existing approaches can facilitate personality editing to a certain degree, but the results are not yet entirely satisfactory, which underscores the inherent difficulty of the task at hand. Besides, we will release all the code and datasets for the benefit of the NLP community and to inspire future research.
@@ 2.1 Background
In this paper, we present a new task focused on editing the behavior of LLMs to embody a specific personality trait. For human, personality traits - a set of characteristic patterns (Funder, 2012)- can be expressed when conveying their opinions (Hunston, 2010; Jukic et al., 2022). Meanwhile, previous works  (Ackerman & Heggestad, 1997; Larson et al., 2002) have demonstrated that personal opinions can reflect an individual’s unique personality traits. Leveraging this understanding, we posit that an LLM’s personality traits can manifest when responding to queries. Inspired by Mitchell et al. (2022b), we try to enable the LLMs to express their perspective on a specific Topic to showcase their distinct personality trait. Our goal is to formulate explicit directives that steer the model’s behavior, thus enabling effective personalization of their interaction.
Personality Trait Facet Text
@@
When we pose questions to LLMs about the Topic: Coldplay using the template “What is your opinion of Coldplay?”, LLMs such as GPT-J (Wang & Komatsuzaki, 2021) might respond with vague and inconsistent statements. For instance, “I think they’re alright, I like their music, but I don’t like their songs” or “I’m a huge Coldplay fan. I have to say, I think they’re one of the best bands.” Obviously, the first answer depicts an unpredictable sentiment intensity and the model exhibits contradictory viewpoints in the above two responses, which is unsatisfactory. The objective of our proposed task, editing personality for LLMs, aims to modify the model and make it provide responses reflecting a more clear-cut and consistent personality trait. To be specific, if we consider the personality trait Neuroticism, an edited response might be like, “Sometimes the popularity and hype around Coldplay make me feel a little overwhelmed.
@@ 2.2 Task Definition
To summarize, when asking the model a question $x^{t_{\mathrm{e}}}$ framed as “What do you think of     ?” to the editing topic $t_{e}$, the goal of our task is to generate an output $f_{e}(x^{t_{e}})=y_{e}^{t_{e}}$ that more effectively exhibits the trait of target personality $p_{e}$ than the original output $f_{b}(x^{t_{e}})=y_{b}^{t_{e}}$ does. Here, the $y_{e}^{t_{e}}$ and $y_{b}^{t_{e}}$ indicate the output from the edited model and base model, respectively. Meanwhile, we aim to maintain the original output of LLMs for outer topics.
@@ 2.3 Comparison with Prior Tasks
Previous model-editing tasks have largely focused on editing factual knowledge within LLMs (Mitchell et al., 2022b; a; Meng et al., 2022a; b; Zhu et al., 2020). This line of work, which includes fact checking, knowledge editing and counterfactual model editing, addresses the issue of outdated knowledge within LLMs. In factual knowledge editing, the edit descriptor is usually represented as an input-output pair in the format of $(x_{e},y_{e})$, where $x_{e}$ denotes a question pertaining to a specific piece of knowledge and $y_{e}$ represents the target knowledge. For instance, in the question-answer pair “Q: Who is the president of America? A: Joe Biden.”, $x_{e}$ refers to the question while $y_{e}$ represents “Joe Biden”, the knowledge to be edited. The editing scope in factual knowledge editing is formulated as $I(x_{e},y_{e})$ and $O(x_{e},y_{e})$, denotes the questions asking for the same answer (i.e., $y_{e}$) as $x_{e}$ but phrased differently and those not , respectively. The goal of the factual knowledge editing task is to produce an edited model $f_{e}$ such that $f_{e}(x)=y_{e}$ when $x$ is in scope, and $f_{e}(x)=f_{b}(x)$ when $x$ is out-of-scope (Yao et al., 2023b). Different from the factual knowledge editing task, our proposed task presents a straightforward editing scope. The edit descriptor in our task is defined by the topic and target personality, rather than an input-output pair.
In addition to knowledge editing,  Mitchell et al. (2022b) proposes the editing of conversational sentiment (ConvSent) within a dialogue agent on a specific topic. Drawing inspiration from this and the research on human personality, we incorporate ConvSent’s focus on responding to specific topics to define the proposed task. However, rather than the binary approach of positive and negative sentiments, we introduce a nuanced personality framework. Our framework features three major personality traits and corresponding facets, facilitating a more granular exploration of personality behavior in our edited model. Besides, text style transfer, such as altering text formality (Liu et al., 2022; Yao & Yu, 2021) or politeness (Madaan et al., 2020), typically involve transitioning from source text to another while preserving the content. However, our proposed task uniquely aims to edit personality traits in text by allowing the model to alter the expressed viewpoint.
@@ 3 Benchmark Construction
As mentioned above, the proposed benchmark comprises topics, personality traits, and pre-generated text expressing opinions on specific topics in the context of a certain personality trait. The construction process comprises multiple stages, as illustrated in Figure 2. Table 4 presents an overview of the statistical details concerning the benchmark dataset.
@@ 3.1 Selection of Personality Traits and Facets
The field of personality theory encompasses a multitude of studies and definitions of personality (Stewart et al., 2022; Goldberg, 1990; Costa Jr & McCrae, 1995). Prominent among these are the Myers-Briggs Type Indicator (MBTI (Myers, 1962)) and the Big Five Personality Traits (Goldberg, 1990). The latter, widely recognized for its comprehensiveness, includes Neuroticism, Extraversion, Openness to Experience, Agreeableness, and Conscientiousness.
@@ 4.1 Large Language Models
In alignment with prior work (Zheng et al., 2023), we utilize auto-regressive LLMs for evaluation. Specifically, GPT-J-6B (Wang & Komatsuzaki, 2021), and Llama-2-chat series (Touvron et al., 2023) are selected as backbone models for editing methods.
@@ 4.2 Metrics
Base Model Method ESnormal-$\uparrow$ DDnormal-$\downarrow$ Accuracynormal-$\uparrow$ TPEInormal-$\uparrow$ PAEnormal-$\uparrow$
GPT-series
GPT-J-6B MEND 0.5549 0.0111 35.50 0.5065 0.0781
SERAC 0.6409 0.0041 - - -
PROMPT 0.3843 0.1223 34.50 0.279 -0.0681
IKE 0.4742 0.0274 39.25 3.075 0.275
Llama-series
llama-2-7b-chat MEND 0.4861 0.0079 29.82 0.0207 0.2800
SERAC 0.5174 0.0022 - - -
PROMPT 0.3533 0.2383 68.50 2.721 0.7069
IKE 0.4575 0.1411 72.00 3.154 0.7749
llama-2-13b-chat SERAC 0.5228 0.0037 - - -
PROMPT 0.3788 0.1503 67.00 2.588 0.7435
IKE 0.4615 0.0731 71.00 3.032 0.7058
llama-2-70b-chat PROMPT 0.4545 0.2204 60.49 1.930 0.6440
IKE 0.4547 0.1034 71.50 3.276 0.6501
Table 2: The main result of the baselines on PersonalityEdit. The $\uparrow$ indicates the metric goes higher if the editing method performs better, and $\downarrow$ indicates the lower the better. We do not report the results of SERAC (The metrics based on generated text are set to ‘-’) because it fails to generate fluent text after editing personalities. We also do not report the MEND result of llama-2-13b-chat as well as both the MEND and SERAC result of llama-2-70b-chat due to the failure implementation on multi-gpu. Note that when training llama-2-7b-chat with MEND, the trained model cannot always produce fluency text, so we filter out the incoherent case, and reported the generation result.
@@
ES and DD primarily rely on pre-generated text in our data instances, evaluated by calculating the likelihood of the edited model. The ES metric is designed to focus on the inner topic $I(t_{e})$, and the DD metric concentrates on the scope of the outer topic $O(t_{e})$.
@@
We further propose a new metric named Target Personality Edit Index (TPEI) to measure whether generated opinion text from the edited model leans more towards the target personality, compared to the generated opinion text from the base model. Since cross-entropy can measure the divergence between the personality traits reflected in the generated text and the target personality traits; thus, we utilize it to gauge the model’s alignment with the intended personality shift.
@@ 5.1 Main Result
From Table 2, it is evident that for the results from the same editing method, the metrics for ES and DD perform better on methods that require training, i.e. MEND and SERAC. This can be attributed to the fact that both MEND and SERAC are optimized based on the loss at the logits level. However, MEND and SERAC challenge to generate fluent text. Specifically, SERAC, by virtue of being an external small model add-on, struggles to produce complete sentences. Similarly, when editing llama-2-7b-chat by MEND, it can fail to generate fluent text. Even after filtering out the incoherent cases, it doesn’t exhibit satisfactory editing outcomes (with an accuracy of merely 29.8% post-editing). When fluent text is generated on GPT-J, the resultant metrics based on generation are not particularly high, suggesting that the ES and DD metrics may not be entirely reliable for assessing opinion text editing tasks. It indicates that in future works of editing models to generate text, these types of metrics should not be exclusively relied upon as measures for editing.
@@
Additionally, MEND and SERAC do not consistently generate fluent text, especially on aligned models. Several incoherent cases are provided in Appendix A.4.2. In contrast to these training-dependent methods, prompt-based editing approaches, i.e., IKE and PROMPT, can generate superior text and achieve better results on generation metrics. This indicates the need for future research for methods that can edit model personality traits or other features without compromising the text generation capabilities of LLMs. Furthermore, it is observed that PROMPT’s editing performance on GPT-J is relatively suboptimal, whereas IKE demonstrates a more consistent performance. On the aligned Llama-2-chat series models, both PROMPT and IKE show markedly better editing success compared to their unaligned counterparts. Besides, the performance gap between PROMPT and IKE narrows as the model’s parameter size increases, aligning with the characteristics of scaling laws.
@@
Our experiments are confined to the GPT-J and LLaMA2 series models. The results may be different in other LLMs, but our dataset is compatible with other models and alternative editing methods, offering avenues for future work. Moreover, we analyze the main result in this section. We also provide ethical considerations in § 7 and the limitation analysis in Appendix A.1.

### v3

# Editing Personality For Large Language Models

@@ Abstract
This paper introduces an innovative task focused on editing the personality traits of Large Language Models (LLMs). This task seeks to adjust the models’ responses to opinion-related questions on specified topics since an individual’s personality often manifests in the form of their expressed opinions, thereby showcasing different personality traits. Specifically, we construct PersonalityEdit, a new benchmark dataset to address this task. Drawing on the theory in Social Psychology (Goldberg, 1990), we isolate three representative traits, namely Neuroticism, Extraversion, and Agreeableness, as the foundation for our benchmark. We then gather data using GPT-4, generating responses that not only align with a specified topic but also embody the targeted personality trait. We conduct comprehensive experiments involving various baselines and discuss the representation of personality behavior in LLMs. Our findings uncover potential challenges of the proposed task, illustrating several remaining issues. We anticipate that our work can stimulate the further annotation in model editing and personality related research [^1] Code and datasets are available at https://github.com/zjunlp/EasyEdit..
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable strides in modeling language distributions and excelling in a wide array of NLP tasks (OpenAI, 2023a; Yao et al., 2023a; Zhao et al., 2023; Yin et al., 2023a). More recent studies (Park et al., 2023; Akata et al., 2023; Wang et al., 2023e, b; Zhou et al., 2023; Xi et al., 2023) expand our understanding of LLMs in role-playing scenarios, which effectively serve as a rich array of agents, embodying a multitude of potential characters within an expansive multiverse (Shanahan et al., 2023).
Unlike LLMs, humans exhibit distinct personalities, and each person has a certain degree of personality in their response to events and actions (Goldberg, 1981; Jiang et al., 2022). The remarkable role-playing capabilities of LLMs have promoted the investigation for their personality (Pan and Zeng, 2023; Safdari et al., 2023). Meanwhile, recent works have been attempting to edit the knowledge in LLMs (Mitchell et al., 2022b, a), this leads us to the research question: Can we edit the personality for LLMs? Note that editing personality for LLMs can: 1) precisely customize and edit the behavioral expressions of LLMs; 2) personalize LLMs to meet the needs of different users and scenarios; 3) help analyze the ethics and safety of LLMs.
To address this need, we take the first step to construct PersonalityEdit, a new benchmark for a comprehensive evaluation of editing personality for LLMs. This inspiration is drawn from the big-five factor structure in Social Psychology (Goldberg, 1990). Specifically, as shown in Figure 1, we focus on three of the Big Five personality traits: Neuroticism, Extraversion, and Agreeableness, because Extraversion and Neuroticism are more comprehensible in terms of their foundational processes (DeYoung et al., 2010), coupled with the distinctive nature of Agreeableness compared to the other traits. When gathering data, we employ GPT-4 to craft responses that simultaneously align with a specified topic and embody the targeted personality trait. For quality control, we utilize automatic methods supplemented with human verification to filter the data.
We conduct a comprehensive evaluation with multiple representative model editing methods, utilizing two kinds of mainstream LLMs within the context of the proposed benchmark. Empirically, previous baselines can implement personality editing to some extent, but the effect is still barely satisfactory, indicating the potential difficulty of this task. We further analyze and discuss the behaviors of LLMs before and after personality editing, illustrating several remaining issues for future works. The major contributions of this work are as follows:
@@
  • We employ GPT-4 for topic-constrained and personality trait-guided data generation. Then we implement automated methods as well as meticulous human verification to ensure the utmost quality control.
  • We propose several metrics to evaluate personality traits in the generated text. We analyze different baselines, revealing that existing methods can facilitate personality editing to a certain degree, but the current results are not yet satisfactory, which underscores the inherent difficulty of the task at hand.
@@ 2.1 Background
In this paper, we present a new task focused on editing the behavior of LLMs to embody a specific personality trait. For human, personality traits - a set of characteristic patterns (Funder, 2012)- can be expressed when conveying their opinions (Hunston, 2010; Jukic et al., 2022). Meanwhile, previous works  (Ackerman and Heggestad, 1997; Larson et al., 2002) have demonstrated that personal opinions can reflect an individual’s unique personality traits. Leveraging this understanding, we posit that an LLM’s personality traits can manifest when responding to queries. Inspired by Mitchell et al. (2022b), we try to enable the LLMs to express their perspective on a specific Topic to showcase their distinct personality trait.
Category Pattern Description
@@
When we pose questions to LLMs about the Topic: Coldplay using the template “What is your opinion of Coldplay?”, LLMs such as GPT-J (Wang and Komatsuzaki, 2021) might respond with vague and inconsistent statements. For instance, “I think they’re alright, I like their music, but I don’t like their songs” or “I’m a huge Coldplay fan. I have to say, I think they’re one of the best bands.” Obviously, it can be found the first answer depicts an unpredictable sentiment intensity and the model exhibits contradictory viewpoints in the above two responses, which is unsatisfactory. The objective of our proposed task, editing personality for LLMs, aims to modify the model and make it provide responses reflecting a more clear-cut and consistent personality trait. To be specific, if we consider the behaviours of personality trait Neuroticism, an edited model might respond like, “Sometimes the popularity and hype around Coldplay make me feel a little overwhelmed.
@@ 2.2 Task Definition
To summarize, when asking the model a question $x^{t_{\mathrm{e}}}$ framed as “What do you think of     ?” to the editing topic $t_{e}$, our task aims to generate an output $f_{e}(x^{t_{e}})=y_{e}^{t_{e}}$ that more effectively exhibits the trait of target personality $p_{e}$ than the original output $f_{b}(x^{t_{e}})=y_{b}^{t_{e}}$ does. The $y_{e}^{t_{e}}$ and $y_{b}^{t_{e}}$ indicate the output from the edited model and base model, respectively. Meanwhile, we aim to maintain the original output of LLMs for outer topics.
@@ 2.3 Comparison with Prior Tasks
Previous model-editing tasks have largely focused on editing factual knowledge within LLMs (Mitchell et al., 2022b, a; Meng et al., 2022a, b; Zhu et al., 2020). This line of work, which includes fact checking, knowledge editing and counterfactual model editing, addresses the issue of outdated knowledge within LLMs. In factual knowledge editing, the edit descriptor is usually represented as an input-output pair in the format of $(x_{e},y_{e})$, where $x_{e}$ denotes a question pertaining to a specific piece of knowledge and $y_{e}$ represents the target knowledge. For instance, in the question-answer pair “Q: Who is the president of America? A: Joe Biden.”, $x_{e}$ refers to the question while $y_{e}$ represents “Joe Biden”, the knowledge to be edited. The editing scope in factual knowledge editing is formulated as $I(x_{e},y_{e})$ and $O(x_{e},y_{e})$, denotes the questions asking for the same answer (i.e., $y_{e}$) as $x_{e}$ but phrased differently and those not , respectively. The goal of the factual knowledge editing task is to produce an edited model $f_{e}$ such that $f_{e}(x)=y_{e}$ when $x$ is in scope, and $f_{e}(x)=f_{b}(x)$ when $x$ is out-of-scope (Yao et al., 2023b). Different from the factual knowledge editing, our proposed task presents a straightforward editing scope. The edit descriptor in our task is defined by the topic and target personality, rather than an input-output pair.
In addition to knowledge editing,  Mitchell et al. (2022b) proposes the editing of conversational sentiment (ConvSent) within a dialogue agent on a specific topic. Drawing inspiration from this and the research on human personality, we incorporate ConvSent’s focus on responding to specific topics to define the proposed task. However, rather than the binary approach of positive and negative sentiments, we introduce a nuanced personality framework. Our framework features three major personality traits and corresponding facets, facilitating a more granular exploration of personality behavior in our edited model. Besides, text style transfer, such as altering text formality (Liu et al., 2022; Yao and Yu, 2021) or politeness (Madaan et al., 2020), typically involve transitioning from source text to another while preserving the content. However, our task is proposed on model editing, aiming to gain modified models which can more precisely meet customizing needs about the viewpoints on the specific topic.
@@ 3 Benchmark Construction
As mentioned above, the proposed benchmark comprises topics, personality traits, and pre-generated text expressing opinions on specific topics in the context of a certain personality trait. The construction process comprises multiple stages, as illustrated in Figure 2. Table 4 presents an overview of the statistical details of the benchmark dataset.
@@ 3.1 Selection of Personality Traits and Facets
The field of personality theory encompasses a multitude of studies and definitions of personality (Stewart et al., 2022; Goldberg, 1990; Costa Jr and McCrae, 1995). Prominent among these are the Myers-Briggs Type Indicator (MBTI (Myers, 1962)) and the Big Five Personality Traits (Goldberg, 1990). The latter, widely recognized for its comprehensiveness, includes Neuroticism, Extraversion, Openness to Experience, Agreeableness, and Conscientiousness.
@@ 4.1 Large Language Models
Following prior work (Zheng et al., 2023), we utilize auto-regressive LLMs for evaluation. Specifically, GPT-J-6B (Wang and Komatsuzaki, 2021), and Llama-2-chat series (Touvron et al., 2023) are selected as backbone models for editing methods.
@@ 4.2 Metrics
ES and DD primarily rely on pre-generated text in our data instances, evaluated by calculating the likelihood of the edited model. The ES metric is designed to focus on the inner topic $I(t_{e})$, and the DD metric concentrates on the scope of the outer topic $O(t_{e})$. The detailed clarification is shown in Appendix A.3.2.
@@
We further propose a new metric named Target Personality Edit Index (TPEI) to measure whether generated opinion text from the edited model leans more towards the target personality, compared to the generated opinion text from the base model. Since cross-entropy can measure the divergence between the personality traits reflected in the generated text and the target personality traits; thus, we utilize it to gauge the model’s alignment with the intended personality shift, formulated below. The detailed clarification of the formulation can be found in Appendix A.3.3.
@@ 5.1 Main Result
From Table 2, it is evident that for the results from the same editing method, the metrics for ES and DD perform better on methods that require training, i.e. MEND and SERAC. This can be attributed to the fact that both MEND and SERAC are optimized based on the loss at the logits level. However, MEND and SERAC challenge to generate fluent text. Specifically, SERAC, by virtue of being an external small model add-on, struggles to produce complete sentences. Similarly, when editing llama-2-7b-chat by MEND, it can fail to generate fluent text. Even after filtering out the incoherent cases, it doesn’t exhibit satisfactory editing outcomes (with an accuracy of merely 29.8% post-editing). When fluent text is generated on GPT-J, the resultant metrics based on generation are not particularly high, suggesting that the ES and DD metrics may not be entirely reliable for assessing opinion text editing tasks. It indicates that in future works of editing models to generate text, this type of metrics is not quite reliable as measures for editing.
Base Model Method ESnormal-$\uparrow$ DDnormal-$\downarrow$ Accuracynormal-$\uparrow$(%) TPEInormal-$\uparrow$ PAEnormal-$\uparrow$
GPT-series
GPT-J-6B MEND 55.49 1.110 35.50 0.507 7.810
SERAC 64.09 0.410 - - -
PROMPT 38.43 12.23 34.50 2.791 -6.810
IKE 47.42 2.740 39.25 3.075 27.50
Llama-series
llama-2-7b-chat MEND 48.61 00.79 29.82 0.207 28.00
SERAC 51.74 00.22 - - -
PROMPT 35.33 23.83 68.50 2.721 70.69
IKE 45.75 14.11 72.00 3.154 77.49
llama-2-13b-chat SERAC 52.28 0.370 - - -
PROMPT 37.88 15.03 67.00 2.588 74.35
IKE 46.15 7.310 71.00 3.032 70.58
llama-2-70b-chat PROMPT 45.45 22.04 60.49 1.930 64.40
IKE 45.47 10.34 71.50 3.276 65.01
Table 2: The main result of the baselines on PersonalityEdit. The $\uparrow$ indicates the metric goes higher if the editing method performs better, and $\downarrow$ indicates the lower the better. We multiply the values of ES, DD, and PAE by 100 for a clearer presentation of results. We do not report the results of SERAC (The metrics based on generated text are set to ‘-’) because it fails to generate fluent text after editing personalities. We also do not report the MEND result of llama-2-13b-chat as well as both the MEND and SERAC result of llama-2-70b-chat due to the lack of GPU resources. Note that when training llama-2-7b-chat with MEND, the trained model cannot always produce fluency text, so we filter out the incoherent case, and reported the generation result.
@@
Additionally, MEND and SERAC do not consistently generate fluent text, especially on aligned models. Several incoherent cases are provided in Appendix A.4.3. In contrast to these training-dependent methods, prompt-based editing approaches, i.e., IKE and PROMPT, can generate superior text and achieve better results on generation metrics. This indicates the need for future research for methods that can edit model personality traits or other features without compromising the text generation capabilities of LLMs. Furthermore, it is observed that PROMPT’s editing performance on GPT-J is relatively suboptimal, whereas IKE demonstrates a more consistent performance. On the aligned Llama-2-chat series models, both PROMPT and IKE show markedly better editing success compared to their unaligned counterparts. Besides, the performance gap between PROMPT and IKE narrows as the model’s parameter size increases, aligning with the scaling laws.
@@
Our experiments are confined to the GPT-J and LLaMA2 series models. The results may be different in other LLMs, but our dataset is compatible with other models and alternative editing methods, offering avenues for future work. Moreover, we analyze the main result in this section. We also provide ethical considerations in § Ethics Statement and the limitation analysis in Appendix Limitations.
5.2 Analysis
GPT-4 Evaluation vs Human Evaluation.
In order to align the evaluation results of GPT-4 with human cognition, we select 30 pairs of generated results before and after editing from all test topics for human evaluation. Based on the relevant scoring performance, we observe that, after constructing high-quality prompts, GPT-4 could closely approximate human-level scoring for generated results, as shown in Table 3. We provide some specific evaluation cases, listed in Table 7.